# An Overview of CV-MDI-QKD

### v1

@@ IV Composable finite-size security
Composable finite-size security of CV-MDI-QKD was initially studied in Refs. Lupo et al. (2018); Papanastasiou et al. (2023). Finite-size means that we consider a finite number of signals exchanged in the protocol, which automatically leads to the use of estimators for evaluating channel parameters. Composability means that there is a security parameter associated with the protocol that quantifies its distance from an ideal implementation. This is called ‘epsilon’ security and combines the failure probabilities of data-processing steps (PE, EC, and PA).
@@
  • Up to an error probability $\epsilon_{\mathrm{ent}}$, Alice and Bob estimate the entropy of the final key.
@@
where the various terms depend on estimators and worst-case estimators of $n_{\mathrm{pe}}=3$ parameters: the link transmissivities, $\tau_{A}$ and $\tau_{B}$, and the total excess noise (Papanastasiou et al., 2023, Eq. (10))
$\Xi=$ $\frac{1}{2}\Big{(}(1-\tau_{A})(\omega_{A}-1)+(1-\tau_{B})(\omega_{B}-1)\Big{)}$
$+g\sqrt{(1-\tau_{A})(1-\tau_{B})}.$ (12)
The mutual information $\widehat{I}_{AB}$ is calculated from the estimators $\widehat{\tau}_{A}$, $\widehat{\tau}_{B}$, and $\widehat{\Xi}$, while Eve’s Holevo information $[I_{E}]_{\text{wc}}$ is computed from the worst-case estimators $\tau_{A}^{\text{wc}}$, $\tau_{B}^{\text{wc}}$, and $\Xi_{\text{wc}}$. In the presence of a stable configuration, these values can be computed over $n_{\mathrm{bks}}m$ points (see Ref. Papanastasiou et al. (2023) for more details on their explicit expressions). In addition, the error reconciliation efficiency is given by
@@ V Network Complexity
V.1 CV MDI-QKD star network
We now consider how CV-MDI-QKD can be extended from the bipartite case to $N$ users. Since MDI-QKD invariably requires the use of a central untrusted relay, it is natural to consider a star network topology in which each of the $N$ users connects to a single central relay. This approach is made possible by means of a generalized Bell detection at the central relay which creates shared correlations that can be post-processed into a secure key common to all parties (‘key conferencing’). Overall, a CV-MDI-QKD network is a specific example of a Gaussian quantum network with untrusted relays, as generally modeled in Ref. Ghalaii et al. (2022).
The structure of the relay for CV-MDI-QKD is depicted in Fig. 5. Each party prepares coherent states |$\ket{\alpha_{i}}$ which are sent through thermal-loss channels to the detector relay. The detector consists of a cascade of $N-1$ beamsplitters of increasing transmissivity. Each output mode on the left of the diagram undergoes homodyne detection in the $q$ quadrature and the final mode at the bottom is homodyned in the $p$ quadrature. The results of the measurements are announced publicly.
Figure 5: Structure of the detector relay for CV-MDI-QKD conferencing. Figure adapted from Ref. Ottaviani et al. (2019)
@@ V.1 CV MDI-QKD star network
Due to the limitations of linear optics, it is important to remark that similar designs for multi-partite MDI-QKD cannot be implemented with discrete variables. A detailed discussion may be found in the supplementary information of Ref. Ottaviani et al. (2019). Finally, it is also worth to mention that a squeezed version of multi-partite CV MDI-QKD can achieve higher rates than the standard version with coherent states.
@@ V.2 CV-MDI-QKD 3-node star network with post-selection
As with the two user case, post-selection techniques may be applied to such a setup to improve the long distance performance of the star network module. In particular, the 3 end user case has been recently investigated Fletcher and Pirandola (2022). In this protocol the 3 parties prepare states of the form |$\ket{\frac{1}{2}(\kappa_{A}\mathbb{Q}_{i}+i\kappa_{i}^{\prime}\mathbb{P}_{i})}$ again having drawn $q_{i}$ and $p_{i}$ from a zero mean Gaussian and encoding the signs as $\kappa_{i}$ and $\kappa_{i}^{\prime}$ and the magnitudes as $\mathbb{Q}_{i}$ and $\mathbb{P}_{i}$. The structure of the detector is the same as the star network described in Section V.1 with one key change: The detector is randomly switched between performing $N-1\leavevmode\nobreak\ q$ homodyne detections and $1\leavevmode\nobreak\ p$ homodyne detection and $N-1\leavevmode\nobreak\ p$ homodyne detections and $1\leavevmode\nobreak\ q$ homodyne detection. This second configuration is shown in the schematic in Fig. 7. The chosen configuration is publicly announced after the detections.
@@
As in Section III.2 the postselected rate may be computed by performing an integral over the region $\Gamma$ where the single point rate $\tilde{R}$ is positive:
[…]
@@ VI Experimental CV-MDI-QKD
CV-MDI-QKD enhances security by eliminating side-channel vulnerabilities on detection. This security benefit comes at the expense of increased implementation complexity compared to one-way CV-QKD. Consequently, despite significant progress in theory, only a few experimental demonstrations of CV-MDI-QKD have been reported.
@@
In 2015, the earliest proof-of-concept demonstration of CV-MDI-QKD was performed using a free-space optical setup, as illustrated in Fig. 9(a) Pirandola et al. (2015). In this experiment, coherent states were generated by modulating a continuous-wave laser at 1064 nm with amplitude and phase modulators driven by Gaussian noise. Phase synchronization between Alice’s and Bob’s signals was achieved using piezoelectric transducers. The CV Bell measurement was implemented using a balanced beam splitter combined with homodyne detectors measuring orthogonal quadratures. The setup achieved a secret key rate of 0.1 bits per relay use under emulated channel losses of 2% for Alice and 60% for Bob, assuming a reconciliation efficiency of 97%. While this work showed the fundamental viability of CV-MDI-QKD, it did not incorporate essential practical features like fiber-optic transmission, independent laser sources at Alice’s and Bob’s stations operating at telecommunication wavelengths, and clock synchronization between the transmitters, making it unsuitable for practical deployment.
@@
Figure 9: Experimental CV-MDI-QKD based on Gaussian-modulated coherent states. (a)  Pirandola et al. (2015) (b)  Tian et al. (2022) (c)  Hajomer et al. (2022) (d) Hajomer et al. (2025). BS: beam splitter; AM: amplitude modulator; AOM: acousto-optic modulator; PM: phase modulator; FM: Faraday mirror; VOA: variable optical attenuator; PBS: polarizing beam splitter; PD: photodetector; PC: polarization controller; FC, fiber collimator; DL: delay line; AWG: arbitrary waveform generator; HWP: Half-wave plate; QWP: quarter-wave plate; BHD: balanced homodyne detector; DAC: Digital-to-analog converter; ABC: Automatic bias controller; CWDM: Coarse Wavelength division multiplexing; FPGA: Field-programmable gate array. Figures adapted from Pirandola et al. (2015)Tian et al. (2022)Hajomer et al. (2022), and Hajomer et al. (2025)
@@
A more practical CV-MDI-QKD implementation was demonstrated in 2022 Tian et al. (2022), as illustrated in Fig. 9(b). This system employed a 10 km standard single-mode fiber (SMF) channel and independent CW lasers operating at the telecom wavelength of 1550 nm. The two lasers were frequency-locked by sending part of Alice’s laser through a fiber channel to Bob’s station, which interfered with a portion of Bob’s laser. The interference result was then fed as an error signal to the phase-locked loop. The system used a 50 ns pulse carved at a repetition rate of 500 kHz. The pulses were split into quantum pulses modulated by cascaded amplitude and phase modulators, pilot pulses used as LO (coming from Bob’s station) and phase-reference pulses (coming from Alice’s station) to compensate for the fast phase drift between the LO and Alice’s signal. The quantum and pilot pulses were time and polarization multiplexed at Alice’s and Bob’s stations using a combination of polarizing beam splitter, delay line, and Faraday mirror before transmitting them through SMF. At the relay, the quantum and the pilot pulses were first demultiplexed. To estimate the fast phase drift, heterodyne detection was performed by tapping off part of the LO pulses and interfering them with the reference pulses in a 90-degree optical hybrid measuring the amplitude quadrature $X_{R}$ and the phase quadrature $P_{R}$. The measurement outcomes were then used to estimate the fast phase drift for each signal pulse as $\Delta\theta=\tan^{-1}(P_{R}/X_{R})$. As for the slow phase drift estimation, Alice and Bob used another set of phase-calibration pulses, which were different from the phase-reference pulses. After estimating the total phase drift, Alice and Bob applied quadrature remapping to their data. With this frequency and phase locking system, a low excess noise of 0.0045 SNU was achieved, enabling a key rate of 0.19 bit per relay use at an error correction efficiency of 97%. This work demonstrated the feasibility of building a CV-MDI-QKD system over fiber channels. However, the complexity of the experiment is rather high. For instance, a separate heterodyne receiver is required to implement the complex frequency and phase locking system, and because it is a pulsed system, an additional amplitude modulator is required for pulse carving. The combination of polarization and time multiplexing also adds another fold of complexity. Moreover, the experiment does not take into account how to synchronize distant transmitters.
@@
The coexistence of CV-MDI-QKD with classical communication channels is pivotal for real-world adoption. A recent system design has addressed this issue, taking into account practical aspects such as clock synchronization and the utilization of SMF channels Hajomer et al. (2025). In this system, an asymmetric network configuration was employed, with the relay being co-located at Alice’s station, as shown in Fig. 9(D). Both senders, Alice and Bob, utilized the same DSP module for quantum state preparation, following the approach in Ref. Hajomer et al. (2022). To achieve synchronization between Alice’s and Bob’s stations, a 10 MHz reference clock was utilized at Bob’s station to generate an optical clock at 1310 nm, which co-propagated with the quantum signal to Alice’s station. Furthermore, real-time phase locking was integrated into the system to lock two independent lasers, eliminating the need for an additional fiber channel as described in Ref. Tian et al. (2022). This significantly simplifies the system structure, making it more suitable for practical adoption. The system operated at a symbol rate of 20 Mbaud and demonstrated key generation in finite-size regimes. This resulted in a two-order-of-magnitude improvement in the key rate compared to previous demonstrations Tian et al. (2022); Pirandola et al. (2015). Despite these advances, further improvements are possible: for instance, enhanced DSP could remove the need for explicit laser phase locking and clock synchronization. Moreover, the achieved symbol rates still lag behind the state-of-the-art in prepare-and-measure coherent-state QKD systems Hajomer et al. (2024) due to bandwidth constraints of the high-efficiency relay and the stringent timing requirements at the relay beam splitter.
@@ VII Conclusions
Acknowledgements. UKRI supported this work through the Integrated Quantum Networks (IQN) Research Hub (EPSRC, Grant No. EP/Z533208/1) and the project ‘QAssure’ (Innovate UK, Grant No. 10102791). AH and TG acknowledge support from the Innovation Foundation Grand Solutions project CyberQ (grant agreement 3200-00035B), the Danish National Research Foundation, Center for Macroscopic Quantum States (bigQ, DNRF142), and the Carlsberg Foundation (grant agreement number CF21-0466). This project has also received funding from the European Union’s Horizon Europe research and innovation programme under the project “Quantum Security Networks Partnership” (QSNP, grant agreement No 101114043).

### v3

@@ IV Composable finite-size security
Composable finite-size security of CV-MDI-QKD was initially studied in Refs. Lupo et al. (2018); Papanastasiou et al. (2023) and later extended to free-space scenarios in Ref. Ghalaii and Pirandola (2023), including satellite implementations. Finite-size means that we consider a finite number of signals exchanged in the protocol, which automatically leads to the use of estimators for evaluating channel parameters. Composability means that there is a security parameter associated with the protocol that quantifies its distance from an ideal implementation. This is called ‘epsilon’ security and combines the failure probabilities of data-processing steps (PE, EC, and PA).
@@
  • Up to an error probability $\epsilon_{\mathrm{ent}}$, Alice and Bob estimate the entropy of the key.
@@
where the various terms depend on estimators and worst-case estimators of $n_{\mathrm{pe}}=3$ parameters: the link transmissivities, $\tau_{A}$ and $\tau_{B}$, and the total excess noise (Papanastasiou et al., 2023, Eq. (10)). For an uncorrelated two-mode attack with two independent entangling cloners ($g^{\prime}=g=0$), the total excess noise takes the form
$\Xi=\frac{1}{2}\Big{(}\tau_{A}\zeta_{A}+\tau_{B}\zeta_{B}\Big{)},$ (12)
where $\zeta_{k}:=(1-\tau_{k})(\omega_{k}-1)/\tau$ is the excess noise of an individual link $k=A$ (Alice-relay) or $B$ (Bob-relay).
The mutual information $\widehat{I}_{AB}$ is calculated from the estimators $\widehat{\tau}_{A}$, $\widehat{\tau}_{B}$, and $\widehat{\Xi}$, while Eve’s Holevo information $I_{E}]_{\text{wc}}$ is computed from the worst-case estimators $\tau_{A}^{\text{wc}}$, $\tau_{B}^{\text{wc}}$, and $\Xi_{\text{wc}}$. In the presence of a stable configuration, these values can be computed over $n_{\mathrm{bks}}m$ points (see Ref. Papanastasiou et al. (2023) for more details on their explicit expressions). In addition, the error reconciliation efficiency is given by
@@
In Fig. 5, we show the composable-secure key rate of the CV-MDI-QKD protocol in the symmetric and asymmetric configurations. We can see how the requirement for composable security takes a toll in terms of achievable distance by comparing the results with those in Fig. 3. For example, in the asymmetric configuration, the asymptotic key rate stretches up to about $100$ km, while this distance is reduced to about $25$ km for the composable key rate.
Figure 5: Composable-secure secret key rate (bits/use) of the symmetric (red line) and asymmetric (black line) CV-MDI-QKD protocol versus fibre distance with standard loss-rate 0.20.2dB/km (the distance in the plot is the sum of the fibre distance of the two links). The distance of Alice from the relay is 55m for the asymmetric configuration. The values of the excess noise for the two individual links are $\zeta_{A}=\zeta_{B}=0.01$ and we assume the two-mode uncorrelated attack ($g^{\prime}=g=0$). We have set $\widehat{\xi}=0.98$, $p_{\text{ec}}=0.95$, $d=14$ and a block size of $N=10^{7}$. The electronic noise and detection efficiency of the relay are set to $u_{\text{el}}=0.01$ and $\eta_{\text{eff}}=0.98$. All epsilon parameters are individually set to $2^{-32}$, while the PE ratio and the modulation variance are optimized. See Ref. Papanastasiou et al. (2023) for details about all the parameters.
@@ V Network Complexity
In this section, we discuss how CV-MDI-QKD can be extended to network scenarios. First, we consider an extension to a star network with an arbitrary number of users. We then consider the specific case of a 3-node star network, for which we introduce the technique of post-selection, with the aim of improving the distribution distance from the central relay.
V.1 CV-MDI-QKD star network
Let us consider how CV-MDI-QKD can be extended from the bipartite case to $N$ users. Since MDI-QKD invariably requires the use of a central untrusted relay, it is natural to consider a star network topology in which each of the $N$ users connects to a single central relay. This approach is made possible by means of a generalized Bell detection at the central relay which creates shared correlations that can be post-processed into a secure key common to all parties (‘key conferencing’). Overall, a CV-MDI-QKD network is a specific example of a Gaussian quantum network with untrusted relays, as generally modeled in Ref. Ghalaii et al. (2022).
[…]
Figure 6: Structure of the detector relay for CV-MDI-QKD conferencing. Figure adapted from Ref. Ottaviani et al. (2019).
@@ V.1 CV-MDI-QKD star network
Due to the limitations of linear optics, it is important to remark that similar designs for multi-partite MDI-QKD cannot be implemented with discrete variables. A detailed discussion may be found in the supplementary information of Ref. Ottaviani et al. (2019). Finally, it is also worth to mention that a squeezed version of multi-partite CV-MDI-QKD can achieve higher rates than the standard version with coherent states. As discussed in Ref. (Ottaviani et al., 2019, Supp. Note 4), higher rates may be achieved if the parties perform the protocol in an entanglement-based configuration where they use TMSV states, sending one mode to the relay while retaining the other. After the generalized Bell detection is performed, they would then apply conditional local squeezing operations to their kept modes before measuring them with heterodyne detection.
@@ V.2 CV-MDI-QKD 3-node star network with post-selection
[…]
@@
As in Section III.2 the postselected rate may be computed by performing an integral over the region $\Gamma$ where the single point rate $\widetilde{R}$ is positive:
$R=\int_{\Gamma}p(\boldsymbol{\mathbb{Q}},\gamma_{q})\ \widetilde{R}(\boldsymbol{\mathbb{Q}},\gamma_{p})\ d\boldsymbol{\mathbb{Q}}\leavevmode\nobreak\ d\gamma_{q}$ (14)
@@ VI Experimental CV-MDI-QKD
CV-MDI-QKD enhances security by eliminating side-channel vulnerabilities on detection. This security benefit comes at the expense of increased implementation complexity compared to one-way CV-QKD. Consequently, despite significant progress in theory, only a few experimental demonstrations of CV-MDI-QKD have been reported, as summarized in Table 1.
@@
In 2015, the earliest proof-of-concept demonstration of CV-MDI-QKD was performed using a free-space optical setup, as illustrated in Fig. 10(a) Pirandola et al. (2015). In this experiment, coherent states were generated by modulating a continuous-wave laser at 1064 nm with amplitude and phase modulators driven by Gaussian noise. Phase synchronization between Alice’s and Bob’s signals was achieved using piezoelectric transducers. The CV Bell measurement was implemented using a balanced beam splitter combined with homodyne detectors measuring orthogonal quadratures. The setup achieved a secret key rate of 0.1 bits per relay use under emulated channel losses of 2% for Alice and 60% for Bob, assuming a reconciliation efficiency of 97%. While this work showed the fundamental viability of CV-MDI-QKD, it did not incorporate essential practical features like fibre-optic transmission, independent laser sources at Alice’s and Bob’s stations operating at telecommunication wavelengths, and clock synchronization between the transmitters, making it unsuitable for practical deployment.
@@
Figure 10: Experimental CV-MDI-QKD based on Gaussian-modulated coherent states. (a)  Pirandola et al. (2015) (b)  Tian et al. (2022) (c)  Hajomer et al. (2022) (d) Hajomer et al. (2025). BS: beam splitter; AM: amplitude modulator; AOM: acousto-optic modulator; PM: phase modulator; FM: Faraday mirror; VOA: variable optical attenuator; PBS: polarizing beam splitter; PD: photodetector; PC: polarization controller; FC, fibre collimator; DL: delay line; AWG: arbitrary waveform generator; HWP: Half-wave plate; QWP: quarter-wave plate; BHD: balanced homodyne detector; DAC: Digital-to-analog converter; ABC: Automatic bias controller; CWDM: Coarse Wavelength division multiplexing; FPGA: Field-programmable gate array. Figures adapted from Refs. Pirandola et al. (2015)Tian et al. (2022)Hajomer et al. (2022), and Hajomer et al. (2025).
Table 1: Comparison of Experimental Parameters in Recent CV-MDI-QKD Demonstrations
Ref. Configuration channel type loss repetition rate key rate, bits/relay use Security
Pirandola et al. (2015) asymmetric Free space 4 dB 100 kHz 0.1 Asymptotic
Tian et al. (2022) asymmetric Fiber 2.02 dB 500 kHz 0.19 Asymptotic
Hajomer et al. (2022) asymmetric VOA 2 dB 5 MHz 0.12 Asymptotic
Hajomer et al. (2025) asymmetric Fiber 2.5 dB 20 MHz 0.13 Finite-size
@@
A more practical CV-MDI-QKD implementation was demonstrated in 2022 Tian et al. (2022), as illustrated in Fig. 10(b). This system employed a 10 km standard single-mode fibre (SMF) channel and independent CW lasers operating at the telecom wavelength of 1550 nm. The two lasers were frequency-locked by sending part of Alice’s laser through a fibre channel to Bob’s station, which interfered with a portion of Bob’s laser. The interference result was then fed as an error signal to the phase-locked loop. The system used a 50 ns pulse carved at a repetition rate of 500 kHz. The pulses were split into quantum pulses modulated by cascaded amplitude and phase modulators, pilot pulses used as LO (coming from Bob’s station) and phase-reference pulses (coming from Alice’s station) to compensate for the fast phase drift between the LO and Alice’s signal. The quantum and pilot pulses were time and polarization multiplexed at Alice’s and Bob’s stations using a combination of polarizing beam splitter, delay line, and Faraday mirror before transmitting them through SMF. At the relay, the quantum and the pilot pulses were first demultiplexed. To estimate the fast phase drift, heterodyne detection was performed by tapping off part of the LO pulses and interfering them with the reference pulses in a 90-degree optical hybrid measuring the amplitude quadrature $X_{R}$ and the phase quadrature $P_{R}$. The measurement outcomes were then used to estimate the fast phase drift for each signal pulse as $\Delta\theta=\tan^{-1}(P_{R}/X_{R})$. As for the slow phase drift estimation, Alice and Bob used another set of phase-calibration pulses, which were different from the phase-reference pulses. After estimating the total phase drift, Alice and Bob applied quadrature remapping to their data. With this frequency and phase locking system, a low excess noise of 0.0045 SNU was achieved, enabling a key rate of 0.19 bit per relay use at an error correction efficiency of 97%. This work demonstrated the feasibility of building a CV-MDI-QKD system over fibre channels. However, the complexity of the experiment is rather high. For instance, a separate heterodyne receiver is required to implement the complex frequency and phase locking system, and because it is a pulsed system, an additional amplitude modulator is required for pulse carving. The combination of polarization and time multiplexing also adds another fold of complexity. Moreover, the experiment does not take into account how to synchronize distant transmitters.
@@
The coexistence of CV-MDI-QKD with classical communication channels is pivotal for real-world adoption. A recent system design has addressed this issue, taking into account practical aspects such as clock synchronization and the utilization of SMF channels Hajomer et al. (2025). In this system, an asymmetric network configuration was employed, with the relay being co-located at Alice’s station, as shown in Fig. 10(D). Both senders, Alice and Bob, utilized the same DSP module for quantum state preparation, following the approach in Ref. Hajomer et al. (2022). To achieve synchronization between Alice’s and Bob’s stations, a 10 MHz reference clock was utilized at Bob’s station to generate an optical clock at 1310 nm, which co-propagated with the quantum signal to Alice’s station. Furthermore, real-time phase locking was integrated into the system to lock two independent lasers, eliminating the need for an additional fibre channel as described in Ref. Tian et al. (2022). This significantly simplifies the system structure, making it more suitable for practical adoption. The system operated at a symbol rate of 20 Mbaud and demonstrated key generation in finite-size regimes. This resulted in a two-order-of-magnitude improvement in the key rate compared to previous demonstrations Tian et al. (2022); Pirandola et al. (2015).
Despite these advances, further improvements are possible. For instance, enhanced DSP techniques could eliminate the need for explicit laser phase locking and clock synchronization. However, the achieved symbol rates in CV-MDI-QKD still lag behind the state-of-the-art in prepare-and-measure coherent-state QKD systems Hajomer et al. (2024). This limitation arises primarily from two factors: (1) bandwidth constraints in the high-efficiency relay, particularly in the readout electronics of balanced homodyne detectors, which often rely on discrete components with parasitic limitations that restrict both speed and noise performance; and (2) the stringent timing requirements at the relay beam splitter. Unlike one-way protocols, CV-MDI-QKD requires precise temporal overlap of optical pulses from independent sources. As the symbol rate increases and pulse durations shorten, synchronization becomes more challenging, directly impacting the interference visibility and system performance.
@@ VII Conclusions
CV-MDI-QKD presents a promising solution for securing high-rate local area networks (LANs) where multiple devices communicate through a central, potentially untrusted relay. The compatibility with standard telecom components and the potential for high key generation rates over short distances make it especially suitable for densely connected LAN environments, such as data centres or enterprise networks, where secure and efficient key exchange is critical.
Acknowledgements. AH and TG acknowledge support from the Innovation Foundation Grand Solutions project CyberQ (grant agreement 3200-00035B), the Danish National Research Foundation, Center for Macroscopic Quantum States (bigQ, DNRF142), and the Carlsberg Foundation (grant agreement number CF21-0466). This project has also received funding from the European Union’s Horizon Europe research and innovation programme under the project “Quantum Security Networks Partnership” (QSNP, grant agreement No. 101114043).